OpenStudy (anonymous):

write equation of the line containing the given point and parelle to the fiven line (6,-5);2x-7y=4

OpenStudy (anonymous):

as the lines are parallel, they have the same slope. 2x-7y=4 2x-4=7y (2/7)x-(4/7)=y so m=2/7 Also, the line contain (6,-5) then: y-(-5) = m (x - 6) y+5= (2/7) (x-6) 7y + 35 = 2x - 12 2x - 7y = 47

OpenStudy (anonymous):

i dont understand what the equation would be then???

OpenStudy (anonymous):

are all the same playing with the algebra. You can write y+5= (2/7) (x-6) or 2x - 7y = 47 is the same